Use the following exchange quotes to help answer problems 18-19. NYSE NASDAQ Bid Ask Bid Ask $149.55$150.45$149.87$ 150.13 $127.88$128.12127.70 S 128.30 CompanyY 18. Currently you have $120.00 in cash and 1000 shares of Stock Y in your brokerage account. You plan to sell your 1000 shares of Stock Y and use the proceeds to purchase as many shares of Stock X as possible. Assume you cannot borrow or buy partial shares of stock; you will also buy or sell at the best price offered. Based on the quotes provided, the cash balance in your brokerage account after you purchase the shares of Stock X is closest to: A. $(60.89) B. $59.11 C. $87.95 D. $89.24 E. $106.10 19. Based on the bid-ask spreads for Stock X & Y the NASDAQǐsmore liquidthan theNYSEwhile Stock y is more liquid than Stock X. Evaluate the underlined words in italics. True or False? Hint: Calculate spread as a percentage of the midpoint of the bid-ask quotes. A. True B. False 20. During lecture we compared Warren Buffets Berkshire Hathaway (BRK.A), which has a market cap of around $480B, to Amazon (AMZN), which has a market cap of around $790B. We concluded that the larger the market cap of a stock the greater the stocks liquidity Evaluate the underlined words in italics. True or False? A. True B. False 21. During lecture we compared the stock market to the bond market. Which of the following statements ( III) is (are) most likely FALSE: I. Stocks are generally more liquid than bonds II. There is more money invested in the US stock market than the US bond market IIl. The fact that bonds are priced relatively infrequently made things worse during the housing crisis A. I E. None of the statements
Answer #1

1. the balance in account is $120, if I sell 1000 shares of Y in NASDAQ, I can get an amount of $ 128.30*1000= $128,300

the total holding amount in Demat account= $128,420

each share price of X= $150.13

the number of shares can be buy= 128,420/150.13= 855 shares

the cost= 855*150.13= $128,361.15

remained amount in account= $58.15

the nearest answer is option B

2. the asset which carries lesser spread have higher liquidity in markets

spread % for X in NYSE= 0.90/150.45= 0.598%

spread % for X in NASDAQ= 0.26/`50.13= 0.173%

hence the spread for X is higher in NYSE

similarly the spread for Y is higher in NASDAQ

Hence the given statement is False

3. Liquidity of a share may depends on many aspects but not only on its firm value. If Amazon is having more units of common stock and having lesser market price per share than BBH, and vice versa. Hence the liquidity depends purely on demand and supply of stock in market and the perception of investors. Hence the given statement is False

4. statement one is False